# Deviation Inequalities for Rényi Divergence Estimators via Variational Expression

### v1

@@ 4.1. Proof of Theorem 1
where $\tilde{\xi}_{r,d,\sigma}$ is given in (36) below, and
$\bar{\xi}_{r,d,\sigma}:=(r^{2}+4r)(1+r\sigma\sqrt{d})\sigma^{-2}.$
To prove Part $(i)$, fix $\mu,\nu\in\mathcal{P}\big{(}\mathbb{B}_{d}(r)\big{)}$. We may assume that $r>0$ since (8) trivially holds for $r=0$. Since $\mathrm{spt}(\mu),\mathrm{spt}(\nu)\subseteq\mathbb{B}_{d}(r)$, $\max_{1\leq i\leq n}\left\|X_{i}\right\|\vee\left\|Y_{i}\right\|\leq r$. Then
@@
$\frac{\mu*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}=\frac{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-X\right\|^{2}}{2\sigma^{2}}}d\mu(x)}{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-Y\right\|^{2}}{2\sigma^{2}}}d\nu(y)}\leq\frac{e^{-\min_{\left\|x^{\prime}\right\|\leq r}\frac{\left\|x-x^{\prime}\right\|^{2}}{2\sigma^{2}}}}{e^{-\max_{\left\|y\right\|\leq r}\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}}\leq\frac{e^{-\frac{(\left\|x\right\|^{2}-2r\left\|x\right\|)}{2\sigma^{2}}}}{e^{-\frac{(\left\|x\right\|^{2}+r^{2}+2\left\|x\right\|r)}{2\sigma^{2}}}}\leq e^{\frac{r^{2}+4r\left\|x\right\|}{2\sigma^{2}}}.$
By similar steps, the same bounds also hold for the ratios $\nu*\varphi_{\sigma}(x)/\mu*\varphi_{\sigma}(x)$ and $\hat{\nu}_{n}*\varphi_{\sigma}(x)/\hat{\mu}_{n}*\varphi_{\sigma}(x)$. Consequently, we have
@@
For applying Lemma 2, define
$\mathcal{F}_{\sigma}(b,q)$ $:=\{f*\varphi_{\sigma}:f\in\mathcal{F}(b,q)\},$ (30a)
$\mathcal{E}_{\sigma,\alpha}(b,q)$ $:=\{e^{\alpha f}*\varphi_{\sigma}:f\in\mathcal{F}(b,q)\},$ (30b)
$F_{b,q}(x)$ $:=0.5b\big{(}1+\left\|x\right\|^{q}\big{)}.$ (30c)
Set $\mathcal{F}_{\sigma}(b):=\mathcal{F}_{\sigma}(b,1)$, $\mathcal{E}_{\sigma}(b):=\mathcal{E}_{\sigma,1}(b,1)$, $F_{b}(x)=0.5b(1+\left\|x\right\|)$, and for $\mathcal{X}\subseteq\mathbb{R}^{d}$
@@
We will also require a bound on the entropy integrals, $J(1,F_{b_{r,\sigma}}*\varphi_{\sigma},\mathcal{F}_{\sigma}(b_{r,\sigma}),\mu)$ and $J(1,e^{F_{b_{r,\sigma}}}*\varphi_{\sigma},\mathcal{E}_{\sigma}(b_{r,\sigma}),\nu)$. To this end, we first obtain a bound on the covering entropy of $\mathcal{F}_{\sigma}(b)$.
@@
where $c_{d,\beta,q}$ is given in (82) below. The same upper bounds in (33a)-(33b) also holds with $\mathcal{F}_{\sigma}(b,q)$ and $F_{b,q}$ replaced by $\mathcal{E}_{\sigma,\alpha}(b,q)$ and $e^{\alpha F_{b,q}}$, respectively. Moreover, the above upper bounds with $q=1$ also applies to all $q\in[0,1]$ up to a universal multiplicative factor, and $c_{d,\frac{d}{2}+1,q}=O\big{(}d^{6d+9}(9e^{2})^{d}\big{)}$ for all $q\in[0,1]$.
The proof of Lemma 3 is given in Section 5.2 and relies on a systematic analysis of the difference of derivatives of a function belonging to $\mathcal{F}_{\sigma}(b,q)$.
Let $\beta=\frac{d}{2}+1$. It follows from (33b) that for all probability measures $\gamma$ such that $\mathrm{spt}(\gamma)\subseteq\mathbb{B}_{d}(r)$, we have
$\log N\Big{(}\epsilon\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\sigma}(b_{r,\sigma}),\left\|\cdot\right\|_{2,\gamma}\Big{)}\lesssim d^{6d+9}(9e^{2})^{d}(r+1)^{d+2}\sigma^{-d}\epsilon^{-\frac{d}{\beta}}.$
@@
$J(1,F_{b_{r,\sigma}}*\varphi_{\sigma},\mathcal{F}_{\sigma}(b_{r,\sigma}),\mu)$ $:=\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\sigma}(b_{r,\sigma}),\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon$
$\lesssim(3ed)^{3(d+2)}(r+1)^{\frac{d}{2}+1}\sigma^{-\frac{d}{2}},$
where we used that $\int_{0}^{1}\epsilon^{-\frac{d}{2\beta}}d\epsilon\lesssim d$ for $\beta=\frac{d}{2}+1$. Similarly,
$J(1,e^{F_{b_{r,\sigma}}}*\varphi_{\sigma},\mathcal{E}_{\sigma}(b_{r,\sigma}),\nu)\lesssim(3ed)^{3(d+2)}(r+1)^{\frac{d}{2}+1}\sigma^{-\frac{d}{2}}.$ (34)
@@
[35, Lemma 1, Lemma 2] If $\mu\in\mathcal{P}\big{(}\mathbb{R}^{d}\big{)}$ is $\sigma^{2}$-sub-Gaussian, then it is $8d\sigma^{2}$-norm sub-Gaussian. In particular, if $X\sim\mu$ and $\mathbb{E}[X]=0$, then $\left(\mathbb{E}[\left\|X\right\|^{p}]\right)^{\frac{1}{p}}\lesssim\sqrt{dp}\sigma$ and $\mathbb{E}\big{[}e^{t\left\|X\right\|}\big{]}\leq e^{4t^{2}d\sigma^{2}}$.
@@
$F_{b_{r,\sigma}}*\varphi_{\sigma}(x)$ $=\frac{1}{(2\pi)^{\frac{d}{2}}\sigma^{d}}\int_{\mathbb{R}^{d}}F_{b_{r,\sigma}}(y)e^{-\frac{\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
$=\frac{r^{2}+4r}{2(2\pi)^{\frac{d}{2}}\sigma^{d+2}}\int_{\mathbb{R}^{d}}(1+\left\|y\right\|)e^{-\frac{\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
@@
$e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}(x)$ $=\frac{1}{(2\pi)^{\frac{d}{2}}\sigma^{d}}\int_{\mathbb{R}^{d}}e^{F_{b_{r,\sigma}}(y)}e^{-\frac{\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
$\leq\frac{e^{\frac{(r^{2}+4r)(1+\left\|x\right\|)}{2\sigma^{2}}}}{(2\pi)^{\frac{d}{2}}\sigma^{d}}\int_{\mathbb{R}^{d}}e^{\frac{(r^{2}+4r)\left\|y-x\right\|-\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
@@
$t_{n,\sigma,\mu,\nu}$
$\lesssim J(1,F_{b_{r,\sigma}}*\varphi_{\sigma},\mathcal{F}_{\sigma}(b_{r,\sigma}),\mu)\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\mu}+J(1,e^{F_{b_{r,\sigma}}}*\varphi_{\sigma},\mathcal{E}_{\sigma}(b_{r,\sigma}),\nu)\left\|e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}\right\|_{2,\nu}$
@@
$\lesssim\tilde{\xi}_{r,d,\sigma}\xi_{r,d,\sigma},$
where
@@
$\log\left(\frac{\hat{\mu}_{n}*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}\right)$ $\leq\frac{\left\|x\right\|\max_{i}\left\|X_{i}\right\|}{\sigma^{2}}+\frac{\mathbb{E}_{\nu}[\left\|Y\right\|^{2}]+2x\cdot\mathbb{E}[Y]}{2\sigma^{2}}$
@@
$\frac{\nu*\varphi_{\sigma}(x)}{\hat{\mu}_{n}*\varphi_{\sigma}(x)}$ $=\frac{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}d\nu(y)}{\frac{1}{n}\sum_{i=1}^{n}e^{-\frac{\left\|x-X_{i}\right\|^{2}}{2\sigma^{2}}}}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{\mathbb{E}_{\nu}\left[e^{-\frac{(\left\|x\right\|^{2}+\left\|Y\right\|^{2}-2x\cdot Y)}{2\sigma^{2}}}\right]}{e^{-\max_{i}\frac{\left\|x\right\|^{2}+\left\|X_{i}\right\|^{2}-2x\cdot X_{i}}{2\sigma^{2}}}}$
$\leq e^{\max_{i}\frac{(\left\|X_{i}\right\|^{2}-2x\cdot X_{i})}{2\sigma^{2}}}\mathbb{E}_{\nu}\left[e^{\frac{x\cdot Y}{\sigma^{2}}}\right]$
@@
We will apply (27) to arrive at the desired conclusion. However, for this purpose, we require the RHS of (39) be bounded almost surely. Since this does not hold for $\mu\in\mathsf{SG}(L)$ in general due to the presence of the term $\max_{i}\left\|X_{i}\right\|^{2}$, we will use a truncation argument by considering the event
@@
Since $X\sim\mu$ satisfies $\left\|X\right\|_{\psi_{p}}\leq L$ for some $p\geq 2$ and $L\geq 1$, we have from [75, Lemma 2.2.2] that
@@
Setting $t=T_{n,p,\kappa,L}:=2^{\frac{1}{p}}Ln^{\kappa}$ for some $\kappa\geq 0$, we obtain from (40) that
@@
$\frac{\mu*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}$ $=\frac{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}d\mu(y)}{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}d\nu(y)}$
$\leq\frac{\mathbb{E}_{\mu}\left[e^{\frac{x\cdot Y}{\sigma^{2}}}\right]}{\mathbb{E}_{\nu}\left[e^{\frac{2x\cdot Y-\left\|Y\right\|^{2}}{2\sigma^{2}}}\right]}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{e^{\frac{\left\|x\right\|^{2}L}{2\sigma^{4}}+\frac{\left\|x\right\|L}{\sigma^{2}}}}{e^{\frac{2x\cdot\mathbb{E}_{\nu}[Y]-\mathbb{E}_{\nu}\left[\left\|Y\right\|^{2}\right]}{2\sigma^{2}}}}$
@@
$\log N\Big{(}\epsilon\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\sigma}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\left\|\cdot\right\|_{2,\gamma}\Big{)}\lesssim d^{6d+9}(9e^{2})^{d}\big{(}T_{n,p,\kappa,L}+1\big{)}^{d+4}\sigma^{-d}\epsilon^{-\frac{d}{\beta}}.$
@@
$:=\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\sigma}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon$
@@
$=\bar{b}_{n,d,p,\kappa,\sigma,L}\left(0.5+\left\|x\right\|^{2}+d\sigma^{2}\right).$
This implies that for $X^{n}\sim\tilde{\mu}^{\otimes n}$ with $\mathrm{spt}(\tilde{\mu})\subseteq\mathbb{B}_{d}(T_{n,p,\kappa,L})$ and any constant $C>0$,
$\mathbb{E}_{\tilde{\mu}}\left[e^{\left(\frac{\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}(X)}{C}\right)^{q}}\right]\leq e^{\frac{\bar{b}_{n,d,p,\kappa,\sigma,L}^{q}\big{(}0.5+d\sigma^{2}+T_{n,p,\kappa,L}^{2}\big{)}^{q}}{C^{q}}}.$
@@
$\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}(X)\right\|_{\psi_{q}}$ $\leq\frac{\bar{b}_{n,d,p,\kappa,\sigma,L}\big{(}0.5+d\sigma^{2}+T_{n,p,\kappa,L}^{2}\big{)}}{(\log 2)^{\frac{1}{q}}}$
$\leq 4\left(\frac{T_{n,p,\kappa,L}+8dL}{\sigma^{2}}\right)^{2}\frac{\big{(}0.5+d\sigma^{2}+T_{n,p,\kappa,L}^{2}\big{)}}{(\log 2)^{\frac{1}{q}}}$
$\lesssim\left(\frac{3}{2}\right)^{\frac{1}{q}}L^{4}d^{3}\sigma^{-4}n^{4\kappa}.$
@@
$J\big{(}1,\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma},\mathcal{F}_{\sigma}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\tilde{\mu}\big{)}\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\tilde{\mu}}$
@@
Substituting the above in (25), we obtain
$O_{n,1,\sigma,\tilde{\mu}}(z)$ $\gtrsim\frac{\sigma^{4}zn^{1-4\kappa}}{L^{4}d^{3}\log(n+1)},$
$R_{n,1,\sigma,\tilde{\mu}}(z)$ $\gtrsim\frac{nz^{2}}{\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\tilde{\mu}}^{2}}\wedge\frac{nz}{\log(n+1)\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}(X)\right\|_{\psi_{1}}}$
$\gtrsim\frac{n^{1-8\kappa}z^{2}\sigma^{8}}{L^{8}d^{6}}\wedge\frac{n^{1-4\kappa}z\sigma^{4}}{L^{4}d^{3}\log(n+1)}.$
Let $0\leq\tau\leq 1$. Set
$\kappa_{d,\tau}^{\star}$ $:=\frac{\tau}{d+12}$ (42a)
$\hat{\xi}_{d,L,\sigma}$ $:=(3ed)^{3(d+2)}(4L)^{\frac{d}{2}+2}\sigma^{-\frac{d}{2}},$ (42b)
$\check{\xi}_{d,L,\sigma}$ $:=L^{4}d^{3}\sigma^{-4}.$ (42c)
@@
$J\big{(}1,\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma},\mathcal{F}_{\sigma}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\tilde{\mu}\big{)}\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\tilde{\mu}}\lesssim\hat{\xi}_{d,L,\sigma}\check{\xi}_{d,L,\sigma}n^{\frac{\tau}{2}},$
$R_{n,1,\sigma,\tilde{\mu}}(z)\gtrsim\frac{n^{\frac{d+12-8\tau}{d+12}}z^{2}}{\check{\xi}_{d,L,\sigma}^{2}}\wedge\frac{n^{\frac{d+12-4\tau}{d+12}}z}{\check{\xi}_{d,L,\sigma}\log(n+1)},$
$O_{n,1,\sigma,\tilde{\mu}}(z)\gtrsim\frac{n^{\frac{d+12-8\tau}{d+12}}z}{\check{\xi}_{d,L,\sigma}\log(n+1)}.$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq 4e^{-\check{\zeta}_{n,d,\tau}(z)},$

### v3

@@ 4.1. Proof of Theorem 1
where $\tilde{\xi}_{r,d,\sigma}$ is given in (33) below, and $\bar{\xi}_{r,d,\sigma}:=(r^{2}+4r)\big{(}1+r\sigma\sqrt{d}\big{)}\sigma^{-2}$.
To prove Theorem 1-$(i)$, fix $\mu,\nu\in\mathcal{P}\big{(}\mathbb{B}_{d}(r)\big{)}$. We may assume that $r>0$ since (8) trivially holds for $r=0$. Since $\mathrm{spt}(\mu),\mathrm{spt}(\nu)\subseteq\mathbb{B}_{d}(r)$, $\max_{1\leq i\leq n}\left\|X_{i}\right\|\vee\left\|Y_{i}\right\|\leq r$. Then
@@
$\frac{\mu*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}=\frac{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-x^{\prime}\right\|^{2}}{2\sigma^{2}}}d\mu(x^{\prime})}{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}d\nu(y)}\leq\frac{e^{-\min_{\left\|x^{\prime}\right\|\leq r}\frac{\left\|x-x^{\prime}\right\|^{2}}{2\sigma^{2}}}}{e^{-\max_{\left\|y\right\|\leq r}\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}}\leq\frac{e^{-\frac{(\left\|x\right\|^{2}-2r\left\|x\right\|)}{2\sigma^{2}}}}{e^{-\frac{(\left\|x\right\|^{2}+r^{2}+2\left\|x\right\|r)}{2\sigma^{2}}}}\leq e^{\frac{r^{2}+4r\left\|x\right\|}{2\sigma^{2}}}.$
By similar steps, the same bounds also hold for $\nu*\varphi_{\sigma}(x)/\mu*\varphi_{\sigma}(x)$ and $\hat{\nu}_{n}*\varphi_{\sigma}(x)/\hat{\mu}_{n}*\varphi_{\sigma}(x)$. Consequently, we have
@@
For $\mathcal{X}\subseteq\mathbb{R}^{d}$, let
@@
Define $F_{b,q}(x):=0.5b\big{(}1+\left\|x\right\|^{q}\big{)}$,
$\mathcal{F}_{\varphi_{\sigma}}(b,q):=\{f*\varphi_{\sigma}:f\in\mathcal{F}(b,q)\}\quad\mbox{ and }\quad\mathcal{E}_{\varphi_{\sigma},\alpha}(b,q):=\{e^{\alpha f}*\varphi_{\sigma}:f\in\mathcal{F}(b,q)\}.$
Set $F_{b}(x)=0.5b(1+\left\|x\right\|)$, $\mathcal{F}_{\varphi_{\sigma}}(b):=\mathcal{F}_{\varphi_{\sigma}}(b,1)$ and $\mathcal{E}_{\varphi_{\sigma}}(b):=\mathcal{E}_{\varphi_{\sigma},1}(b,1)$.
For applying Lemma 2, we will require a bound on the entropy integrals, $J(1,F_{b_{r,\sigma}}*\varphi_{\sigma},\mathcal{F}_{\varphi_{\sigma}}(b_{r,\sigma}),\mu)$ and $J(1,e^{F_{b_{r,\sigma}}}*\varphi_{\sigma},\mathcal{E}_{\varphi_{\sigma}}(b_{r,\sigma}),\nu)$. To this end, we first obtain a slightly more general result bounding the covering entropy of $\mathcal{F}_{\varphi_{\sigma}}(b,q)$.
@@
where $c_{d,\beta,q}$ is given in (76) below. The same upper bounds in (32a)-(32b) also holds with $\mathcal{F}_{\varphi_{\sigma}}(b,q)$ and $F_{b,q}$ replaced by $\mathcal{E}_{\varphi_{\sigma},\alpha}(b,q)$ and $e^{\alpha F_{b,q}}$, respectively. Moreover, the above upper bounds with $q=1$ also applies to all $q\in[0,1]$ up to a universal multiplicative factor, and $c_{d,\frac{d}{2}+1,q}=O\big{(}d^{6d+9}(9e^{2})^{d}\big{)}$ for all $q\in[0,1]$.
The proof of Lemma 3 is given in Section 5.2 and relies on a systematic analysis of the difference of derivatives of a function belonging to $\mathcal{F}_{\varphi_{\sigma}}(b,q)$.
Let $\beta=\frac{d}{2}+1$. It follows from (32b) that for all probability measures $\gamma$ such that $\mathrm{spt}(\gamma)\subseteq\mathbb{B}_{d}(r)$,
$\log N\Big{(}\epsilon\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\varphi_{\sigma}}(b_{r,\sigma}),\left\|\cdot\right\|_{2,\gamma}\Big{)}\lesssim d^{6d+9}(9e^{2})^{d}(r+1)^{d+2}\sigma^{-d}\epsilon^{-\frac{d}{\beta}}.$
@@
$J(1,F_{b_{r,\sigma}}*\varphi_{\sigma},\mathcal{F}_{\varphi_{\sigma}}(b_{r,\sigma}),\mu)$ $:=\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\varphi_{\sigma}}(b_{r,\sigma}),\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon\lesssim\tilde{\xi}_{r,d,\sigma},$
where we used that $\int_{0}^{1}\epsilon^{-\frac{d}{2\beta}}d\epsilon\lesssim d$ for $\beta=\frac{d}{2}+1$, and
Similarly,
$J(1,e^{F_{b_{r,\sigma}}}*\varphi_{\sigma},\mathcal{E}_{\varphi_{\sigma}}(b_{r,\sigma}),\nu)\lesssim\tilde{\xi}_{r,d,\sigma}.$ (34)
@@
[41, Lemma 1, Lemma 2] If $\mu\in\mathcal{P}\big{(}\mathbb{R}^{d}\big{)}$ is $\sigma^{2}$-sub-Gaussian, then it is $8d\sigma^{2}$-norm sub-Gaussian. In particular, if $X\sim\mu$ and $\mathbb{E}[X]=0$, then $\left(\mathbb{E}[\left\|X\right\|^{p}]\right)^{\frac{1}{p}}\lesssim\sqrt{dp}\sigma$ and $\mathbb{E}\big{[}e^{t\left\|X\right\|}\big{]}\leq e^{4t^{2}d\sigma^{2}}$.
@@
$F_{b_{r,\sigma}}*\varphi_{\sigma}(x)=\frac{1}{(2\pi)^{\frac{d}{2}}\sigma^{d}}\int_{\mathbb{R}^{d}}F_{b_{r,\sigma}}(y)e^{-\frac{\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$ $=\frac{r^{2}+4r}{2(2\pi)^{\frac{d}{2}}\sigma^{d+2}}\int_{\mathbb{R}^{d}}(1+\left\|y\right\|)e^{-\frac{\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
@@
$e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}(x)=\frac{1}{(2\pi)^{\frac{d}{2}}\sigma^{d}}\int_{\mathbb{R}^{d}}e^{F_{b_{r,\sigma}}(y)}e^{-\frac{\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$ $\leq\frac{e^{\frac{(r^{2}+4r)(1+\left\|x\right\|)}{2\sigma^{2}}}}{(2\pi)^{\frac{d}{2}}\sigma^{d}}\int_{\mathbb{R}^{d}}e^{\frac{(r^{2}+4r)\left\|y-x\right\|-\left\|y-x\right\|^{2}}{2\sigma^{2}}}dy$
@@
$t_{n,\varphi_{\sigma},\mu,\nu}$
$\lesssim J(1,F_{b_{r,\sigma}}*\varphi_{\sigma},\mathcal{F}_{\varphi_{\sigma}}(b_{r,\sigma}),\mu)\left\|F_{b_{r,\sigma}}*\varphi_{\sigma}\right\|_{2,\mu}+J(1,e^{F_{b_{r,\sigma}}}*\varphi_{\sigma},\mathcal{E}_{\varphi_{\sigma}}(b_{r,\sigma}),\nu)\left\|e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}\right\|_{2,\nu}$
@@
$\lesssim\tilde{\xi}_{r,d,\sigma}\xi_{r,d,\sigma}.$
@@
$\log\left(\frac{\hat{\mu}_{n}*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}\right)$ $\leq\frac{\left\|x\right\|\max_{i}\left\|X_{i}\right\|}{\sigma^{2}}+\frac{\mathbb{E}_{\nu}[\left\|Y\right\|^{2}]+2x\cdot\mathbb{E}_{\nu}[Y]}{2\sigma^{2}}$
@@
$\frac{\nu*\varphi_{\sigma}(x)}{\hat{\mu}_{n}*\varphi_{\sigma}(x)}=\frac{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}d\nu(y)}{\frac{1}{n}\sum_{i=1}^{n}e^{-\frac{\left\|x-X_{i}\right\|^{2}}{2\sigma^{2}}}}\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{\mathbb{E}_{\nu}\left[e^{-\frac{(\left\|x\right\|^{2}+\left\|Y\right\|^{2}-2x\cdot Y)}{2\sigma^{2}}}\right]}{e^{-\max_{i}\frac{\left\|x\right\|^{2}+\left\|X_{i}\right\|^{2}-2x\cdot X_{i}}{2\sigma^{2}}}}$ $\leq e^{\max_{i}\frac{(\left\|X_{i}\right\|^{2}-2x\cdot X_{i})}{2\sigma^{2}}}\mathbb{E}_{\nu}\left[e^{\frac{x\cdot Y}{\sigma^{2}}}\right]$
@@
We will apply (28) to arrive at the desired conclusion. However, for this purpose, we require the RHS of (38) be bounded a.s. Since this does not hold for $\mu\in\mathsf{SG}(L)$ in general due to the presence of the term $\max_{i}\left\|X_{i}\right\|^{2}$, we will use a truncation argument by considering the event
@@
Since $X\sim\mu$ satisfies $\left\|X\right\|_{\psi_{p}}\leq L$ for some $p\geq 2$ and $L\geq 1$, we have from [83, Lemma 2.2.2] (a prefactor of $K=2^{\frac{1}{p}}$ suffices) that
@@
Setting
$t=T_{n,p,\kappa,L}:=2^{\frac{1}{p}}Ln^{\kappa},$
for some $\kappa\geq 0$, we obtain from (39) that
@@
$\frac{\mu*\varphi_{\sigma}(x)}{\nu*\varphi_{\sigma}(x)}=\frac{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}d\mu(y)}{\int_{\mathbb{R}^{d}}e^{-\frac{\left\|x-y\right\|^{2}}{2\sigma^{2}}}d\nu(y)}\leq\frac{\mathbb{E}_{\mu}\left[e^{\frac{x\cdot Y}{\sigma^{2}}}\right]}{\mathbb{E}_{\nu}\left[e^{\frac{2x\cdot Y-\left\|Y\right\|^{2}}{2\sigma^{2}}}\right]}$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{e^{\frac{\left\|x\right\|^{2}L}{2\sigma^{4}}+\frac{\left\|x\right\|L}{\sigma^{2}}}}{e^{\frac{2x\cdot\mathbb{E}_{\nu}[Y]-\mathbb{E}_{\nu}\left[\left\|Y\right\|^{2}\right]}{2\sigma^{2}}}}$
@@
$\log N\Big{(}\epsilon\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\varphi_{\sigma}}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\left\|\cdot\right\|_{2,\gamma}\Big{)}\lesssim d^{6d+9}(9e^{2})^{d}\big{(}T_{n,p,\kappa,L}+1\big{)}^{d+4}\sigma^{-d}\epsilon^{-\frac{d}{\beta}}.$
@@
$:=\int_{0}^{1}\sup_{\gamma}\sqrt{\log N\Big{(}\epsilon\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\gamma},\mathcal{F}_{\varphi_{\sigma}}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\left\|\cdot\right\|_{2,\gamma}\Big{)}}d\epsilon$
@@
$\lesssim\bar{b}_{n,d,p,\kappa,\sigma,L}\left(0.5+\left\|x\right\|^{2}+d\sigma^{2}\right).$
@@
$\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{\infty,\tilde{\mu}}\leq\bar{b}_{n,d,p,\kappa,\sigma,L}\left(0.5+T_{n,p,\kappa,L}^{2}+d\sigma^{2}\right)\lesssim L^{4}d^{3}\sigma^{-4}n^{4\kappa}.$
@@
$J\big{(}1,\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma},\mathcal{F}_{\varphi_{\sigma}}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\tilde{\mu}\big{)}\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\tilde{\mu}}$
@@
Let $0\leq\tau\leq 1$. Set $\kappa_{d,\tau}^{\star}:=\frac{\tau}{d+12}$,
$\hat{\xi}_{d,L,\sigma}$ $:=(3ed)^{3(d+2)}(4L)^{\frac{d}{2}+2}\sigma^{-\frac{d}{2}}\quad\mbox{and}\quad\check{\xi}_{d,L,\sigma}:=L^{4}d^{3}\sigma^{-4}.$ (41)
@@
$J\big{(}1,\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma},\mathcal{F}_{\varphi_{\sigma}}\big{(}\bar{b}_{n,d,p,\kappa,\sigma,L},2\big{)},\tilde{\mu}\big{)}\left\|\bar{F}_{n,d,p,\kappa,\sigma,L}*\varphi_{\sigma}\right\|_{2,\tilde{\mu}}\lesssim\hat{\xi}_{d,L,\sigma}\check{\xi}_{d,L,\sigma}n^{\frac{\tau}{2}}.$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq e^{-\check{\zeta}_{n,d,\tau}(z)},$